### v2

@@ 1 Introduction
Predicting the neutron distribution in space, energy, and time is essential when modeling inertial confinement fusion experiments, pulsed neutron sources, and nuclear criticality safety experiments, among other systems. The behavior of neutrons can be modeled with a Monte Carlo simulation, where particles are created and transported to produce a particle history [2]. The path of a particle and the specific set of events that occur within its history are governed by pseudorandom numbers, known probabilities (e.g., from material data), and known geometries. Data about how particles move and/or interact within the system are tallied to compute quantities of interest with an associated statistical error from the Monte Carlo process.
@@
Two methods are commonly used to sample the random walk in a Monte Carlo neutron transport algorithm: surface tracking [2] and Woodcock delta tracking [3], which we simply call “delta tracking” for the remainder of this work. These two tracking algorithms have complementary performance bottlenecks [4]. In geometrically complex models with many surfaces, surface tracking can require many expensive calculations to find the distance to a nearest surface which delta tracking avoids. In contrast, when delta tracking in models with certain material compositions, the required rejection sampling can dominate computational cost. Furthermore, standard implementations of delta tracking preclude evaluating quantities of interest with a track-length estimator, instead opting for a collision estimator, while surface tracking has no such restrictions [5]. The track-length estimator often provides lower-variance estimates of quantities of interest (see Section 2) [6]. Due to these differences, relative performance of these methods depend on the problem. For a certain class of problems, a hybrid method may allow greater performance than either approach used individually. There have been other recent developments on delta tracking—namely weighted delta tracking algorithms [7, 4]—but in this work we explore variants of the non-weighted version.
@@
Previous research into hybrid delta-surface tracking on a structured mesh showed good performance for problems with complex arrangements of optically thin materials [8]. However, this work was limited to problems that define material regions and track on a structured mesh. Making material-based decisions about when to use delta and surface tracking has also been explored and implemented in production Monte Carlo codes [9, 10, 11]. Here, we extend the idea of tracking on a structured mesh to full delta tracking (eliminating the distance-to-surface check almost entirely) and tallying to a structured mesh, allowing the relatively efficient use of a track-length estimator for scalar flux. We describe, verify, and evaluate the performance of a delta-tracking algorithm that allows the use of a track-length estimator on a structured tally mesh.
@@
We then implement this delta tracking plus track length estimator combination in two hybrid delta-tracking schemes: one in which the choice of tracking algorithm is based on the material region, and another based on the particle energy. We implement this work in Monte Carlo / Dynamic Code (MC/DC), an open-source Monte Carlo neutron transport application designed to conduct rapid numerical methods development, specifically for time-dependent problems [12]. We compute figures of merit for four computationally difficult revOnetime-dependent fixed source benchmark problems and run-time results for a whole CPU node (two Intel x86 Xeon Sapphire Rapids CPUs) and a whole GPU node (four Nvidia Tesla V100 GPUs) of supercomputing systems at Lawrence Livermore National Laboratory.
@@
This work demonstrates the first published use of a track-length estimator for scalar flux with full delta tracking, the first use of delta tracking in conjunction with continuously moving surfaces, and the first use of a hybrid delta-tracking method based on particle energy. The methods we present here will not be used to compute reaction rate tallies. While this does limit the general applicability of these methods, scalar flux is a necessary quantity for hybrid methods (e.g., iterative quasi Monte Carlo [13, 14]) and other uses.
@@ 2 Tracking Algorithms and Estimators
Conducting Monte Carlo calculations in problems with heterogeneous materials requires some method of treating discontinuities in the cumulative probability distribution function [15]. This is done with a tracking (or sampling) method. The first method we consider is surface tracking, described in Algorithm 1. For a particle in material $m$, the distance to collision is sampled from a cumulative probability distribution function by
@@
This sampling of the cumulative probability distribution function will only hold true while the particle is in material $m$. If the distance to collision is beyond a material interface in a system with multiple materials, the particle must be stopped at that interface surface and a new distance to collision is sampled using the new material’s $\Sigma_{t,m}(E)$. This approach is an unbiased way to deal with the sampling of the distance to collision in a heterogeneous medium. In a standard surface-tracking algorithm, both the distance to collision ($d_{\text{collision}}$) and the distance to the nearest surface along the particle’s direction of travel ($d_{\text{surface}}$) are computed, and the smaller of these two distances determines which event happens to the particle: a collision or a surface crossing. After or while the particle is moving, tallies can be accumulated to compute quantities of interest. If a collision occurs, more sampling and associated operations are performed (e.g., isotropic scattering of a particle off a particular constituent nuclide). If the particle is still alive and has not exited the problem domain, the algorithm is repeated. The computation of the distance to the nearest surface can become quite expensive as geometries grow in complexity (e.g., complex combinatorial solid geometries (CSG) or CAD-based surfaces). Surface tracking is at the heart of many modern Monte Carlo neutron transport applications, including MCNP [16], Shift [17, 18, 19], GREAPMC [20], MONK/MCBEND [11], MCATK [21], Mercury [22], McCARD [23, 24], SCONE [25], COG [26], TRIPOLI-4/5 [27, 28], and OpenMC [29].
@@
An added complication when using delta tracking is that only certain tallies can be efficiently scored. Many estimators can be used to compute quantities of interest (e.g., scalar flux, reaction rates) by tallying events that occur within a given region of phase space. Two common scalar flux estimators are the collision estimator and the track-length (also known as the path length), estimator [2]. Often, tallies are scored into bins on a structured mesh that overlays the surfaces and material regions that a Monte Carlo simulation uses to conduct actual transport operations. The track-length estimator is
@@
The collision estimator will often produce a flux estimate with higher variance compared to the track-length estimator when $\Sigma_{t,m}$ is small (i.e., optically thin, less-dense materials) and is characterized by a zero tally in a true void region [6, 5]. On the other hand, the track-length estimator will tally into every mesh bin the particle moves through during its random walk. In some problem regimes, more information will be scored for a given number of particle histories, which will result in a tally with lower variance.
The estimators in Eqs. (5) and  (6) only apply to flux integrals. Response functions for other reaction rates (e.g., fission rate density) will require knowledge of the macroscopic cross section of that reaction in a given location [15]. In this work, when using the track-length estimator with delta tracking we limit ourselves to flux integrals only. We discuss this further in Section 7.
Delta-tracking algorithms are implemented in many production Monte Carlo neutron transport applications, including Serpent2 [10, 5, 9, 30], MONK/MCBEND [11], SCONE [25], IMPC [31], TRIPOLI-5 [28], and GUARDYAN [32]. Notably, the MONK Monte Carlo neutron transport code is the direct successor to the GEM code where Woodcock et al. first implemented delta tracking [3].
Modern transport applications often involve only one of these two tracking algorithms, and the implementation of their code base is optimized from there. However, either method of sampling the probability distribution functions is valid for the same system in the same simulation, even at the same location in phase space. Some developers have taken advantage of this logic. For example, Serpent2 Monte Carlo code establishes regions of delta tracking and surface tracking based on the ratio between $\Sigma_{t,m}$ and $\Sigma_{\scriptstyle\text{maj}}$ and a user-supplied constant $\in[0,1]$ (where 0 means delta tracking only and $1$ surface tracking only) [9]. MONK/MCBEND supports material regions inside of which delta tracking is implemented (called hole geometries) for complex materials like stochastically heterogeneous materials (e.g., TRISO fuel elements, aggregate and cement mixtures in concrete) while surface tracking is used elsewhere [11, 33].
@@ 3 Hybrid Delta Tracking Schemes
In this section, we introduce and verify the voxelized tally structure that allows the implementation in MC/DC of a track-length estimator with delta tracking. We also introduce two hybrid surface-delta tracking schemes that we implemented for time-dependent transport with moving surfaces in MC/DC. The first is a region-based delta tracking method we call “hybrid-in-material” that has been previously explored in other production Monte Carlo neutron transport applications [31, 9]. The second method uses delta tracking only for neutrons with energies above resolved neutron cross section resonances, where total cross sections are often similar to the majorant, and mean free paths are long relative to nominal system dimensions. This approach is called “hybrid-in-energy”.
@@ 3.1 Voxelized Tallies
Standard delta-tracking algorithms do not require information about the material or physical mesh cell a given particle occupies at any moment in transport. Conventional wisdom dictates that repeatedly determining locations, cross sections, and tally bin indices while using delta tracking would be costly and eliminate any benefit to performance that the method provides [6]. However, there is nothing mathematically preventing the use of a track-length estimator with delta tracking. Furthermore, if all that is desired as a quantity of interest for a given problem is the scalar flux, no cross section lookups are needed while tallying—only the physical track length and corresponding bin. This forms the underlying idea behind the voxelized tally method. We investigate whether efficient implementation of tallying allows for effective use of a track length estimator when delta tracking and if this results in a performance benefit.
@@
As of v0.11.1, MC/DC [12, 34] has a tally algorithm that scores track lengths crossing multiple structured mesh tally bins (voxels) in a single operation. This algorithm is based on a sweep method in which the initial state [mesh cell index, ($x$, $y$, $z$) position, direction of travel, speed, and particle clock] is known, as is the distance to the next event. The particle is then swept from tally voxel to tally voxel, accumulating the track length traveled in a given voxel along the way. Figure 2 shows a hypothetical particle track and the voxels to be tallied (in red) in a single operation. In MC/DC’s surface-tracking algorithm, this occurs even before moving a particle to an event. This is computationally efficient as we are tallying to a voxel—a structured rectilinear mesh bin.
@@
To verify that MC/DC’s voxelized track-length tallies with delta tracking converge to the correct solution and at the correct rate, we use four analytic fixed-source benchmark problems from MC/DC’s verification suite. revOneA fixed-source calculation is well posed when the location of the source is known and an initial condition is proved (i.e. not $\alpha$ or $k$ eigenvalue calculations). We compare the error ($\epsilon$) in both the $L_{\infty}$ and $L_{2}$ norms, from integral quantities of interest to a reference solution and plot the error as a function of increasing particle count. The convergence rate should be the standard Monte Carlo convergence rate of $N^{-1/2}$. Our verification problems are:
  • AZURV1 time-dependent benchmark both super and sub critical (Figure 3) [35];
  • A time-dependent infinite pin cell using the 371 group SHEM cross sections (Figure 4) [36];
  • Reed’s problem (Figure 5) revTwo[37]; and
@@
Figure 3: Convergence rate verification of AZURV1 [35], showing the expect Monte Carlo convergence rate ($N^{-1/2}$).
@@
Figure 4: Convergence rate of an infinite pin using the SHEM 361 group cross section library [36], showing the expected Monte Carlo convergence rate ($N^{-1/2}$).
Figure 5: Convergence rate of flux from Reed’s problem [37] showing the expect Monte Carlo convergence rate ($N^{-1/2}$).
@@ 3.2 Hybrid-in-Material
The first hybrid method we implement in MC/DC is the hybrid-in-material method. Each active particle is assigned an additional flag that declares which transport algorithm it is using to sample a distance from the next event. This is similar to Serpent2’s algorithm [5]; however, our voxelized tally structure allows us to use track-length estimators in all regions, not just those in which the particle is undergoing surface tracking. Also, Serpent2’s algorithm makes automatic decisions about where to surface and delta track based on a single user-supplied cut-off value between zero and one for the whole problem [9], whereas we leave it the user to specify on a material-by-material basis. A more automatic approach similar to Serpent2 will be implemented in future work. The IMPC Monte Carlo neutron transport code also implements a similar hybrid-in-material surface-delta tracking scheme [31].
@@ 3.3 Hybrid-in-Energy
Most high-energy neutron interactions with materials are characterized by relatively long mean free paths (i.e., small cross sections) because they primarily undergo potential scattering [38]. In systems with a large number of surfaces and fast neutrons, delta tracking generally does better compared to surface tracking, because surface tracking gets stuck moving particles from region to region. Delta tracking, on the other hand, will stream particles through the whole problem, never conducting expensive distance-to-nearest-surface computations. Furthermore, at higher energies, cross sections tend to vary less across material interfaces, alleviating issues with the rejection-sampling loop. These observations motivate the development of the “hybrid-in-energy” method that employs delta tracking for high energy neutrons and surface tracking, otherwise.
For example, consider a continuous energy version of the C5G7 benchmark reactor geometry [39]. revOneDue to the lack of isotopes available in MC/DC’s cross section library revOneTo simplify the problem to allow for rapid numerical methods exploration we replace the guide tube and fission chamber materials with moderator material. We also replace MOX fuels with UO2 with different enrichment. The material composition is given by Table LABEL:tab:c5ce in Appendix B. Figure 1 shows the macroscopic cross sections for the seven-material reactor, with the majorant in black. Around $10\text{\,}\mathrm{keV}$ the neutron resonances end and the system is “high energy”.

### v3

@@ 1 Introduction
Predicting the neutron distribution in space, energy, and time is essential when modeling inertial confinement fusion experiments, pulsed neutron sources, and nuclear criticality safety experiments, among other systems. The behavior of neutrons can be modeled with a Monte Carlo simulation, where particles are created and transported to produce a particle history [Lewis1984]. The path of a particle and the specific set of events that occur within its history are governed by pseudorandom numbers, known probabilities (e.g., from material data), and known geometries. Data about how particles move and/or interact within the system are tallied to compute quantities of interest with an associated statistical error from the Monte Carlo process.
@@
Two methods are commonly used to sample the random walk in a Monte Carlo neutron transport algorithm: surface tracking [Lewis1984] and Woodcock delta tracking [woodcock_1965_deltatracking], which we simply call “delta tracking” for the remainder of this work. These two tracking algorithms have complementary performance bottlenecks [morgan_weighted-delta-tracking_2015]. In geometrically complex models with many surfaces, surface tracking can require many expensive calculations to find the distance to a nearest surface which delta tracking avoids. In contrast, when delta tracking in models with certain material compositions, the required rejection sampling can dominate computational cost. Furthermore, standard implementations of delta tracking preclude evaluating quantities of interest with a track-length estimator, instead opting for a collision estimator, while surface tracking has no such restrictions [leppanen_2017_collision]. The track-length estimator often provides lower-variance estimates of quantities of interest (see Section 2) [mc2018]. Due to these differences, relative performance of these methods depend on the problem. For a certain class of problems, a hybrid method may allow greater performance than either approach used individually. There have been other recent developments on delta tracking—namely weighted delta tracking algorithms [molnar_variance_2018, morgan_weighted-delta-tracking_2015]—but in this work we explore variants of the non-weighted version.
@@
Previous research into hybrid delta-surface tracking on a structured mesh showed good performance for problems with complex arrangements of optically thin materials [morgan_2021_mcatk]. However, this work was limited to problems that define material regions and track on a structured mesh. Making material-based decisions about when to use delta and surface tracking has also been explored and implemented in production Monte Carlo codes [leppanen_development_2013, leppanen_2010_burnup, richards_monk_2015]. Here, we extend the idea of tracking on a structured mesh to full delta tracking (eliminating the distance-to-surface check almost entirely) and tallying to a structured mesh, allowing the relatively efficient use of a track-length estimator for scalar flux. We describe, verify, and evaluate the performance of a delta-tracking algorithm that allows the use of a track-length estimator on a structured tally mesh.
@@
We then implement this delta tracking plus track length estimator combination in two hybrid delta-tracking schemes: one in which the choice of tracking algorithm is based on the material region, and another based on the particle energy. We implement this work in Monte Carlo / Dynamic Code (MC/DC), an open-source Monte Carlo neutron transport application designed to conduct rapid numerical methods development, specifically for time-dependent problems [morgan_monte_2024]. We compute figures of merit for four computationally difficult time-dependent fixed source benchmark problems and run-time results for a whole CPU node (two Intel x86 Xeon Sapphire Rapids CPUs) and a whole GPU node (four Nvidia Tesla V100 GPUs) of supercomputing systems at Lawrence Livermore National Laboratory.
@@
This work demonstrates the first published use of a track-length estimator for scalar flux with full delta tracking, the first use of delta tracking in conjunction with continuously moving surfaces, and the first use of a hybrid delta-tracking method based on particle energy. The methods we present here will not be used to compute reaction rate tallies. While this does limit the general applicability of these methods, scalar flux is a necessary quantity for hybrid methods (e.g., iterative quasi Monte Carlo [Pasmann03062023, Pasmann04042025]) and other uses.
@@ 2 Tracking Algorithms and Estimators
Conducting Monte Carlo calculations in problems with heterogeneous materials requires some method of treating discontinuities in the cumulative probability distribution function [lux_1998]. This is done with a tracking (or sampling) method. The first method we consider is surface tracking, described in Algorithm 1. For a particle in material $m$, the distance to collision is sampled from a cumulative probability distribution function by
@@
This sampling of the cumulative probability distribution function will only hold true while the particle is in material $m$. If the distance to collision is beyond a material interface in a system with multiple materials, the particle must be stopped at that interface surface and a new distance to collision is sampled using the new material’s $\Sigma_{t,m}(E)$. This approach is an unbiased way to deal with the sampling of the distance to collision in a heterogeneous medium. In a standard surface-tracking algorithm, both the distance to collision ($d_{\text{collision}}$) and the distance to the nearest surface along the particle’s direction of travel ($d_{\text{surface}}$) are computed, and the smaller of these two distances determines which event happens to the particle: a collision or a surface crossing. After or while the particle is moving, tallies can be accumulated to compute quantities of interest. If a collision occurs, more sampling and associated operations are performed (e.g., isotropic scattering of a particle off a particular constituent nuclide). If the particle is still alive and has not exited the problem domain, the algorithm is repeated. The computation of the distance to the nearest surface can become quite expensive as geometries grow in complexity (e.g., complex combinatorial solid geometries (CSG) or CAD-based surfaces). Surface tracking is at the heart of many modern Monte Carlo neutron transport applications, including MCNP [MCNP_RisingArmstrongEtAl], Shift [hamilton_continuous-energy_2019, pandya_implementation_2016, shift2], GREAPMC [gearpmc], MONK/MCBEND [richards_monk_2015], MCATK [mcatk], Mercury [POZULP2024], McCARD [mccard1, mccard2], SCONE [scone], COG [cog], TRIPOLI-4/5 [trip4, trip5], and OpenMC [romano_openmc_2015].
@@
An added complication when using delta tracking is that only certain tallies can be efficiently scored. Many estimators can be used to compute quantities of interest (e.g., scalar flux, reaction rates) by tallying events that occur within a given region of phase space. Two common scalar flux estimators are the collision estimator and the track-length (also known as the path length), estimator [Lewis1984]. Often, tallies are scored into bins on a structured mesh that overlays the surfaces and material regions that a Monte Carlo simulation uses to conduct actual transport operations. The track-length estimator is
@@
The collision estimator will often produce a flux estimate with higher variance compared to the track-length estimator when $\Sigma_{t,m}$ is small (i.e., optically thin, less-dense materials) and is characterized by a zero tally in a true void region [mc2018, leppanen_2017_collision]. On the other hand, the track-length estimator will tally into every mesh bin the particle moves through during its random walk. In some problem regimes, more information will be scored for a given number of particle histories, which will result in a tally with lower variance.
The estimators in Eqs. (5) and  (6) only apply to flux integrals. Response functions for other reaction rates (e.g., fission rate density) will require knowledge of the macroscopic cross section of that reaction in a given location [lux_1998]. In this work, when using the track-length estimator with delta tracking we limit ourselves to flux integrals only. We discuss this further in Section 7.
Delta-tracking algorithms are implemented in many production Monte Carlo neutron transport applications, including Serpent2 [leppanen_2010_burnup, leppanen_2017_collision, leppanen_development_2013, leppanen_2015_serpent], MONK/MCBEND [richards_monk_2015], SCONE [scone], IMPC [fang_development_2022], TRIPOLI-5 [trip5], and GUARDYAN [molnar_guardyan_2019]. Notably, the MONK Monte Carlo neutron transport code is the direct successor to the GEM code where Woodcock et al. first implemented delta tracking [woodcock_1965_deltatracking].
Modern transport applications often involve only one of these two tracking algorithms, and the implementation of their code base is optimized from there. However, either method of sampling the probability distribution functions is valid for the same system in the same simulation, even at the same location in phase space. Some developers have taken advantage of this logic. For example, Serpent2 Monte Carlo code establishes regions of delta tracking and surface tracking based on the ratio between $\Sigma_{t,m}$ and $\Sigma_{\scriptstyle\text{maj}}$ and a user-supplied constant $\in[0,1]$ (where 0 means delta tracking only and $1$ surface tracking only) [leppanen_development_2013]. MONK/MCBEND supports material regions inside of which delta tracking is implemented (called hole geometries) for complex materials like stochastically heterogeneous materials (e.g., TRISO fuel elements, aggregate and cement mixtures in concrete) while surface tracking is used elsewhere [richards_monk_2015, richards_monk_2025].
@@ 3 Hybrid Delta Tracking Schemes
In this section, we introduce and verify the voxelized tally structure that allows the implementation in MC/DC of a track-length estimator with delta tracking. We also introduce two hybrid surface-delta tracking schemes that we implemented for time-dependent transport with moving surfaces in MC/DC. The first is a region-based delta tracking method we call “hybrid-in-material” that has been previously explored in other production Monte Carlo neutron transport applications [fang_development_2022, leppanen_development_2013]. The second method uses delta tracking only for neutrons with energies above resolved neutron cross section resonances, where total cross sections are often similar to the majorant, and mean free paths are long relative to nominal system dimensions. This approach is called “hybrid-in-energy”.
@@ 3.1 Voxelized Tallies
Standard delta-tracking algorithms do not require information about the material or physical mesh cell a given particle occupies at any moment in transport. Conventional wisdom dictates that repeatedly determining locations, cross sections, and tally bin indices while using delta tracking would be costly and eliminate any benefit to performance that the method provides [mc2018]. However, there is nothing mathematically preventing the use of a track-length estimator with delta tracking. Furthermore, if all that is desired as a quantity of interest for a given problem is the scalar flux, no cross section lookups are needed while tallying—only the physical track length and corresponding bin. This forms the underlying idea behind the voxelized tally method. We investigate whether efficient implementation of tallying allows for effective use of a track length estimator when delta tracking and if this results in a performance benefit.
@@
As of v0.11.1, MC/DC [morgan_monte_2024, transport_cement_mcdc_2024] has a tally algorithm that scores track lengths crossing multiple structured mesh tally bins (voxels) in a single operation. This algorithm is based on a sweep method in which the initial state [mesh cell index, ($x$, $y$, $z$) position, direction of travel, speed, and particle clock] is known, as is the distance to the next event. The particle is then swept from tally voxel to tally voxel, accumulating the track length traveled in a given voxel along the way. Figure 2 shows a hypothetical particle track and the voxels to be tallied (in red) in a single operation. In MC/DC’s surface-tracking algorithm, this occurs even before moving a particle to an event. This is computationally efficient as we are tallying to a voxel—a structured rectilinear mesh bin.
@@
To verify that MC/DC’s voxelized track-length tallies with delta tracking converge to the correct solution and at the correct rate, we use four analytic fixed-source benchmark problems from MC/DC’s verification suite. A fixed-source calculation is well posed when the location of the source is known and an initial condition is proved (i.e. not $\alpha$ or $k$ eigenvalue calculations). We compare the error ($\epsilon$) in both the $L_{\infty}$ and $L_{2}$ norms, from integral quantities of interest to a reference solution and plot the error as a function of increasing particle count. The convergence rate should be the standard Monte Carlo convergence rate of $N^{-1/2}$. Our verification problems are:
  • AZURV1 time-dependent benchmark both super and sub critical (Figure 3) [ganapol_2001_homogeneous];
  • A time-dependent infinite pin cell using the 371 group SHEM cross sections (Figure 4) [hfaiedh_2005_shem];
  • Reed’s problem (Figure 5) [reed_difference_1971]; and
@@
Figure 3: Convergence rate verification of AZURV1 [ganapol_2001_homogeneous], showing the expect Monte Carlo convergence rate ($N^{-1/2}$).
@@
Figure 4: Convergence rate of an infinite pin using the SHEM 361 group cross section library [hfaiedh_2005_shem], showing the expected Monte Carlo convergence rate ($N^{-1/2}$).
Figure 5: Convergence rate of flux from Reed’s problem [reed_difference_1971] showing the expect Monte Carlo convergence rate ($N^{-1/2}$).
@@ 3.2 Hybrid-in-Material
The first hybrid method we implement in MC/DC is the hybrid-in-material method. Each active particle is assigned an additional flag that declares which transport algorithm it is using to sample a distance from the next event. This is similar to Serpent2’s algorithm [leppanen_2017_collision]; however, our voxelized tally structure allows us to use track-length estimators in all regions, not just those in which the particle is undergoing surface tracking. Also, Serpent2’s algorithm makes automatic decisions about where to surface and delta track based on a single user-supplied cut-off value between zero and one for the whole problem [leppanen_development_2013], whereas we leave it the user to specify on a material-by-material basis. A more automatic approach similar to Serpent2 will be implemented in future work. The IMPC Monte Carlo neutron transport code also implements a similar hybrid-in-material surface-delta tracking scheme [fang_development_2022].
@@ 3.3 Hybrid-in-Energy
Most high-energy neutron interactions with materials are characterized by relatively long mean free paths (i.e., small cross sections) because they primarily undergo potential scattering [duderstadt_hamilton]. In systems with a large number of surfaces and fast neutrons, delta tracking generally does better compared to surface tracking, because surface tracking gets stuck moving particles from region to region. Delta tracking, on the other hand, will stream particles through the whole problem, never conducting expensive distance-to-nearest-surface computations. Furthermore, at higher energies, cross sections tend to vary less across material interfaces, alleviating issues with the rejection-sampling loop. These observations motivate the development of the “hybrid-in-energy” method that employs delta tracking for high energy neutrons and surface tracking, otherwise.
For example, consider a continuous energy version of the C5G7 benchmark reactor geometry [hou2017oecd]. Due to the lack of isotopes available in MC/DC’s cross section library To simplify the problem to allow for rapid numerical methods exploration we replace the guide tube and fission chamber materials with moderator material. We also replace MOX fuels with UO2 with different enrichment. The material composition is given by Table LABEL:tab:c5ce in Appendix B. Figure 1 shows the macroscopic cross sections for the seven-material reactor, with the majorant in black. Around $10\text{\,}\mathrm{keV}$ the neutron resonances end and the system is “high energy”.